A man bought a cow and a calf for $360, paying eight times as much for the cow
as for the calf. What was the cost of each?

Answer by Maths68(1474)


Let
Price of the calf = x
Price of the cow = 8x
Total amount paid = 360
Price of calf + Price of cow = Total amount paid
x+8x=360
9x=360
9x/9=360/9
x=40
so

Price of the calf = x = 40
Price of the cow = 8x = 8*40 = 320
Total Amount paid = 40+320=360